OpenStudy (m):

x is exponential random variable f(x) = (1/θ)e^-x/θ given: P(x < 2.40) = .520 find: P(x > 4.36)

OpenStudy (anonymous):

You need to integrate f(x) from 0 to 2.40 and set that equal to .520. You will have an equation you can solve for theta. Once you find theta integrate the pdf again from 4.36 to infinity, or take one minus the integration from 0 to 4.36 whichever you prefer.

OpenStudy (anonymous):

sorry i dont know why it posted that many times

OpenStudy (m):

i keep getting a negative number for the theta -1.5694 i end up with a negative answer which i don't think it's correct

OpenStudy (m):

make that theta = -3.67

OpenStudy (m):

-e^2.40/theta - [-e^-0/theta] = .520 -e^-2.40/theta + 1 = .520 e^2.40/theta = .48 -2.40/theta = ln(.48) theta = 3.2699 i was trying to ln the limit 0 at first and that's why i was having problem. yay for perseverance :)

OpenStudy (m):

-e^-2.40/theta - [-e^-0/theta] = .520 -e^-2.40/theta + 1 = .520 e^-2.40/theta = .48 -2.40/theta = ln(.48) theta = 3.2699 -e-x/theta from 4.36 to infinity -e^-inf/3.269 - [-e^-4.36/3.269] = 0 + .2635 =.2635

OpenStudy (anonymous):

yep I think that's correct